# Testing the Predictions of Surprisal Theory in 11 Languages

### v3

@@ 3.2 Language Models
One recent study has hypothesized that, when deriving surprisal estimates for psycholinguistic modeling, the size of the context window can bias estimates (Hoover et al., 2022). Their reasoning is that short context windows could shift probability mass away from very low-frequency words, which would be better predicted from longer contexts. Therefore, we estimate surprisal and contextual entropy from mGPT in two contexts: In short contexts the model is given only the current sentence (up until the current word); in long contexts we use the model’s full input window size of 512 characters. We use long contexts for our first analysis, and use both contexts for our second analysis, which investigates both the shape of the reading times–surprisal linking function and the influence of context length on these results.
@@ 3.3 Regression Models
All of our regression models are fit to predict the reading time $y(w_{t},\boldsymbol{w}_{<t})$ of a word $w_{t}$ in a context $\boldsymbol{w}_{<t}$ from the predictor vector $\mathbf{x}_{t}$. In addition to looking at the word $w_{t}$, our predictor includes quantities derived from the previous two words $w_{t-1},w_{t-2}$ to control for potential spillover effects. We will refer to the three words $w_{t},w_{t-1},w_{t-2}$ as our regressor words. Following previous work in this area, all regression models include the word length and log-unigram frequency, as estimated by Speer (2022), for all regressor words in a predictor $\mathbf{x}_{t}$ for a specific index $t$. The predictors above constitute our (context invariant) baseline predictors. Regression models are trained and evaluated using 10-fold cross validation. For more information on the regressions used in each of our experiments, see Appendix A. The significance of the observed $\Delta$ values between target and baseline models is assessed via a paired permutation test that checks whether $\Delta$ is significantly different from zero. We use permutation tests because our comparisons because they make no assumption about the distribution of the test statistic. Instead, the test uses the empirical distribution of differences in likelihoods, as estimated using averages computed over permutations of likelihoods, in order to compute $p$-values.
@@ 4.1 Surprisal
For an aggregate test of the effects of surprisal, we fit an additional regression model on the combined data from all languages to predict gaze duration with random by-language effects. We use a fully maximal random effect structure, as advocated in Barr et al. (2013). We find that the model with surprisal leads to significantly greater than zero $\Delta$ in all cases ($p<0.001$). Although surprisal leads to a positive $\Delta$ across languages, we do observe some variation in the magnitude of this effect, or the predictive power obtained by regression model. For both mGPT and monoT(all) we observe the highest predictive power in Russian and Dutch, with lower predictive power in Spanish, English, and Hebrew. One natural question to ask is whether imbalances in the model’s training data leads to some of this variation—do models make better predictions for language where they have seen more data? However, there are converging pieces of evidence from our data suggesting that differences in dataset size is not the main cause of the by-language variation. First, both mGPT and monoT(all) show relatively lower predictive power for some large-data languages such as Spanish and English. Second, and quite interestingly, similar patterns of predictive power can be observed for our monoT(30m) models, where training dataset size is controlled across languages. Here, as with the other models, we observe larger values of $\Delta$ in Dutch and Russian and smaller values of $\Delta$ in English, Spanish and Hebrew. These results pose a puzzle, as the languages for which the models obtain higher $\Delta$ are not obviously different from those for which the models obtain lower $\Delta$, in terms of their linguistic features. For example, English (lower $\Delta$) and Dutch (higher $\Delta$) are both Western Germanic. Further investigation is needed to determine if these patterns hold up for other crosslinguistic reading time datasets.
@@ 4.2 Contextual Entropy
Results can be seen in Figure 2, where the replace regression is indicated with a triangle and the add regression is indicated with a circle. First, we find that replacing surprisal with entropy tends to hurt predictive power in most cases. For example, for mGPT, $\Delta$ is negative in 6/11 languages and significantly so in two (Dutch ($p<0.05$) and Italian ($p<0.05$)), implying overfitting. Negative effects are even stronger for the monoT(all) model, where we find negative gaze duration $\Delta$ in every language (results are significant in 5/11). Adding entropy as an additional predictor, on the other hand, generally improves the model’s predictive power. For example, for mGPT and gaze duration, $\Delta$ from the add regression is positive in 8/11 languages, and significantly so in 5 (English, Greek, Korean, Russian and Turkish). In addition, $\Delta$ is significantly positive for the add regression for all three reading time measures when data is combined across languages, as shown in the ‘All’ column at the left of Figure 2. Results are less strong for monoT(all), where positive $\Delta$ shows up predominantly for first fixation. As before, we run an aggregate test with data from all languages including by-language random effects. [^7] Following the same methodology as the previous test, we look at the effect of adding or replacing surprisal across all regressor words. For gaze duration, we find that adding contextual entropy leads to positive $\Delta$ (mGPT, $p<0.001$; monoT(all), $p<0.01$) and that replacement leads to negative $\Delta$ (mGPT, $p<0.01$; monoT(all), $p<0.001$). Overall, we take these results as being in line with those reported in Pimentel et al. (2023). Our findings suggest that contextual entropy has a weak—albeit consistent—effect on reading times across languages, and therefore that participants may be pre-planning their processing times based on the expected surprisal of upcoming words.
@@ 4.3 Variation Across Languages
The correlations can be seen in Figure 4. We do find a relatively strong negative correlation across languages, however it is not significant ($\rho=-0.497,p=0.1$). We do not find any evidence of correlation in the language family data. Although the negative by-language correlation suggests that, for languages where mGPT has lower perplexity, it may be a better model of psycholinguistic behavior, the lack of significance is in line with the negative results from Japanese.
@@ 4.4 Model Coefficients
Turning to contextual entropy, we find slightly smaller effects, and slightly more variance between languages. There is no obvious relationship between the effect sizes for surprisal and contextual entropy. For example, Dutch, which has a larger surprisal effect, has one of the smallest effect sizes for entropy. For frequency, we find a consistently negative effect for $w_{t}$, as expected—as words get more frequent they take less time to read. For $w_{t-1}$ and $w_{t-2}$ effects are much smaller and less consistent across languages. For example, Dutch, Finnish, Italian and Russian all have consistently positive frequency effects for $w_{t-1}$, whereas in Turkish and Greek, these effects are negative.
@@ 6.3 Concurrent Work
We want to briefly note the differences between the work presented here and a concurrent study that also used the MECO dataset (i.e., de Varda and Marelli, 2022). While de Varda and Marelli’s research questions are similar to ours, their methods and conclusions are quite different. Instead of an autoregressive language model, they use a masked language model (mBERT; Devlin et al., 2019), which has access to both left and right context. An issue with this strategy is that the surprisal values produced by this setup are not psychologically plausible estimates of actual surprisals, which are estimated from the left context alone.  [^12] Because the perceptual span is limited to about 14 characters to the right of a fixation (Rayner, 1975) and little linguistic information is gleaned from the far right of the perceptual span (Schotter et al., 2012), upcoming word identities cannot have a substantial causal influence on a word’s first-pass reading behavior (Granger, 1969). which weakens the ability to test psycholinguistic causal claim about the relationship between surprisal and reading times. In their experiments, de Varda and Marelli do not find significant effects of pseudo-surprisal on gaze duration in four of the 12 languages in MECO, [^13] They include Estonian, which we drop as it was not in mGPT’s training data. including English, and find significant effects of pseudo-surprisal on other eye movement measures in even fewer of the languages, which they view as evidence that surprisal might not be a consistent predictor of reading times across languages. [^14] Their study does not consider contextual entropy. While we are aligned on the importance of de Varda and Marelli’s research questions, we believe that their failure to replicate surprisal effects for English—or to find it for other languages—reflects the limitations in their methodological choices.
@@ Appendix A Regression Modeling Details
We give more details on the regression formulae used in the various experiments reported in the main section. Our notation is as follows: reading_time is the reading time of the word of interest, i.e., $w_{t}$, surp is the surprisal of $w_{t}$, prev_surp is the surprisal of the previous word, i.e., $w_{t-1}$, and prev2_surp is the surprisal of the word two previous, i.e., $w_{t-2}$. The other variables use the same prev and prev2 prefixes, and we simply explain the variable names for the current index $t$ for the sake of brevity, below. For these, ent indicates the contextual entropy of $W_{t}$, len indicates the length of $w_{t}$ in characters, and freq indicates the log unigram frequency of $w_{t}$.
@@
For both tests, the baseline model included surprisal, length and unigram frequency, i.e., it was the first model given in the paragraph above. For the replace test, the target regression model we use is
@@ Appendix B Surprisal versus RT for ${w_{t-1}}$
The results are consistent across models used, and suggest that the relationship between reading time and surprisal of the previous word is somewhat variable across languages. For English, Italian, Korean, Russian and Spanish we find a relationship that is roughly linear and increasing, i.e., similar to the results for surprisal of the current word. For Dutch, Turkish and Hebrew, we find a relationship that is roughly increasing, but visually non-linear. For Finnish, German and Greek, we find either a flat or negative relationship. These results are in line with the effect terms plotted in Figure 3, where we find very weak and sometimes negative coefficients for the $w_{t-1}$ surprisal term for these languages (i.e., the middle x-tick position in the top row). Overall, these results are consistent with the linear effect that has been previously observed in English. However, they suggest that the impact of the surprisal of the previous word varies between languages.

### v4

@@ 3.2 Language Models
One recent study has hypothesized that, when deriving surprisal estimates for psycholinguistic modeling, the size of the context window can bias estimates (Hoover et al., 2022). Their reasoning is that short context windows could shift probability mass away from very low-frequency words, which would be better predicted from longer contexts. Therefore, we estimate surprisal and contextual entropy from mGPT in two contexts: In short contexts the model is given only the current sentence (up until the current word); in long contexts we use the model’s full input window size of 512 tokens. We use long contexts for our first analysis, and use both contexts for our second analysis, which investigates both the shape of the reading times–surprisal linking function and the influence of context length on these results.
@@ 3.3 Regression Models
All of our regression models are fit to predict the reading time $y(w_{t},\boldsymbol{w}_{<t})$ of a word $w_{t}$ in a context $\boldsymbol{w}_{<t}$ from the predictor vector $\mathbf{x}_{t}$. In addition to looking at the word $w_{t}$, our predictor includes quantities derived from the previous two words $w_{t-1},w_{t-2}$ to control for potential spillover effects. We will refer to the three words $w_{t},w_{t-1},w_{t-2}$ as our regressor words. Following previous work in this area, all regression models include the word length and negative log frequency (i.e., unigram surprisal), as estimated by Speer (2022), for all regressor words in a predictor $\mathbf{x}_{t}$ for a specific index $t$. The predictors above constitute our (context invariant) baseline predictors. Regression models are trained and evaluated using 10-fold cross validation. For more information on the regressions used in each of our experiments, see Appendix A. The significance of the observed $\Delta$ values between target and baseline models is assessed via a paired permutation test that checks whether $\Delta$ is significantly different from zero. We use permutation tests because our comparisons because they make no assumption about the distribution of the test statistic. Instead, the test uses the empirical distribution of differences in likelihoods, as estimated using averages computed over permutations of likelihoods, in order to compute $p$-values.
@@ 4.1 Surprisal
For an aggregate test of the effects of surprisal, we fit an additional regression model on the combined data from all languages to predict gaze duration with random by-language effects. We use a fully maximal random effect structure, as advocated in Barr et al. (2013). We find that the model with surprisal leads to significantly greater than zero $\Delta$ in all cases ($p<0.001$). Although surprisal leads to a positive $\Delta$ across languages, we do observe some variation in the magnitude of this effect, or the predictive power obtained by regression model. For both mGPT and monoT(all) we observe the highest predictive power in Russian and Dutch, with lower predictive power in Spanish, English, and Hebrew. One natural question to ask is whether imbalances in the model’s training data leads to some of this variation—do models make better predictions for language where they have seen more data? However, there are converging pieces of evidence from our data suggesting that differences in dataset size is not the main cause of the by-language variation. First, both mGPT and monoT(all) show relatively lower predictive power for some large-data languages such as Spanish. Second, and quite interestingly, similar patterns of predictive power can be observed for our monoT(30m) models, where training dataset size is controlled across languages. Here, as with the other models, we observe larger values of $\Delta$ in Italian and Russian and smaller values of $\Delta$ in Spanish and Hebrew. These results pose a puzzle, as the languages for which the models obtain higher $\Delta$ are not obviously different from those for which the models obtain lower $\Delta$, in terms of their linguistic features. For example, Spanish (lower $\Delta$) and Italian (higher $\Delta$) are both Romance languages. Further investigation is needed to determine if these patterns hold up for other crosslinguistic reading time datasets.
@@ 4.2 Contextual Entropy
Results can be seen in Figure 2, where the replace regression is indicated with a triangle and the add regression is indicated with a circle. First, we find that replacing surprisal with entropy tends to hurt predictive power in most cases. For example, for mGPT, gaze duration $\Delta$ is negative in 7/11 languages and significantly so in three (Spanish $p<0.05$, Dutch ($p<0.01$) and Italian ($p<0.05$)), implying overfitting. Negative effects are even stronger for the monoT(all) model, where we find negative gaze duration $\Delta$ in every language except Hebrew (results are significant in 3/11). Adding entropy as an additional predictor, on the other hand, generally improves the model’s predictive power. For example, for mGPT and gaze duration, $\Delta$ from the add regression is positive in 8/11 languages, and significantly so in 3 (English, Greek and Turkish). In addition, $\Delta$ is significantly positive for the add regression for all three reading time measures when data is combined across languages, as shown in the ‘All’ column at the left of Figure 2. Results are less strong for monoT(all), where positive $\Delta$ shows up predominantly for first fixation. As before, we run an aggregate test with data from all languages including by-language random effects. [^7] Following the same methodology as the previous test, we look at the effect of adding or replacing surprisal across all regressor words. For gaze duration, we find that adding contextual entropy leads to positive $\Delta$ (mGPT, $p<0.001$; monoT(all), $p<0.01$) and that replacement leads to negative $\Delta$ (mGPT, $p<0.01$; monoT(all), $p<0.001$). Overall, we take these results as being in line with those reported in Pimentel et al. (2023). Our findings suggest that contextual entropy has a weak—albeit consistent—effect on reading times across languages, and therefore that participants may be pre-planning their processing times based on the expected surprisal of upcoming words.
@@ 4.3 Variation Across Languages
The correlations can be seen in Figure 4. We do find a negative correlation across languages, however it is not significant ($\rho=-0.186,p=0.6$). We do not find any evidence of correlation in the language family data. Although the negative by-language correlation suggests that, for languages where mGPT has lower perplexity, it may be a better model of psycholinguistic behavior, the lack of significance is in line with the negative results from Japanese.
@@ 4.4 Model Coefficients
Turning to contextual entropy, we find slightly smaller effects, and slightly more variance between languages. There is no obvious relationship between the effect sizes for surprisal and contextual entropy. For example, Dutch, which has a larger surprisal effect, has one of the smallest effect sizes for entropy. For negative log frequency (i.e., unigram surprisal), we find a consistently positive effect for $w_{t}$, as expected implying that as words get more frequent they take less time to read. For $w_{t-1}$ and $w_{t-2}$ effects are much smaller and less consistent across languages. For example, Dutch, Finnish, Italian and Russian all have consistently negative frequency effects for $w_{t-1}$, whereas in Turkish and Hebrew, these effects are positive.
@@ 6.3 Concurrent Work
We want to briefly note the differences between the work presented here and a concurrent study that also used the MECO dataset (i.e., de Varda and Marelli, 2022). While de Varda and Marelli’s research questions are similar to ours, their methods and conclusions are quite different. Instead of an autoregressive language model, they use a masked language model (mBERT; Devlin et al., 2019), which has access to both left and right context. An issue with this strategy is that the surprisal values produced by this setup are not psychologically plausible estimates of actual surprisals, which are estimated from the left context alone, [^12] Because the perceptual span is limited to about 14 characters to the right of a fixation (Rayner, 1975) and little linguistic information is gleaned from the far right of the perceptual span (Schotter et al., 2012), upcoming word identities cannot have a substantial causal influence on a word’s first-pass reading behavior (Granger, 1969). which weakens the ability to test psycholinguistic causal claim about the relationship between surprisal and reading times. In their experiments, de Varda and Marelli do not find significant effects of pseudo-surprisal on gaze duration in four of the 12 languages in MECO, [^13] They include Estonian, which we drop as it was not in mGPT’s training data. including English, and find significant effects of pseudo-surprisal on other eye movement measures in even fewer of the languages, which they view as evidence that surprisal might not be a consistent predictor of reading times across languages. [^14] Their study does not consider contextual entropy. While we are aligned on the importance of de Varda and Marelli’s research questions, we believe that their failure to replicate surprisal effects for English—or to find it for other languages—reflects the limitations in their methodological choices.
@@ Appendix A Regression Modeling Details
We give more details on the regression formulae used in the various experiments reported in the main section. Our notation is as follows: reading_time is the reading time of the word of interest, i.e., $w_{t}$, surp is the surprisal of $w_{t}$, prev_surp is the surprisal of the previous word, i.e., $w_{t-1}$, and prev2_surp is the surprisal of the word two previous, i.e., $w_{t-2}$. The other variables use the same prev and prev2 prefixes, and we simply explain the variable names for the current index $t$ for the sake of brevity, below. For these, ent indicates the contextual entropy of $W_{t}$, len indicates the length of $w_{t}$ in characters, and freq indicates the negative log frequency of $w_{t}$.
@@
For both tests, the baseline model included surprisal, length and negative log frequency, i.e., it was the first model given in the paragraph above. For the replace test, the target regression model we use is
@@ Appendix B Surprisal versus RT for ${w_{t-1}}$
The results are consistent across models used, and suggest that the relationship between reading time and surprisal of the previous word is somewhat variable across languages. For English, Italian, Korean, Russian, Greek and Spanish we find a relationship that is roughly linear and increasing, i.e., similar to the results for surprisal of the current word. For Dutch and Turkish, we find a relationship that is roughly increasing, but visually non-linear. For Finnish, German and Hebrew, we find either a negative relationship, or a relationship that is difficult to interpret. These results are in line with the effect terms plotted in Figure 3, where we find very weak and sometimes negative coefficients for the $w_{t-1}$ surprisal term for these languages (i.e., the middle $x$-tick position in the top row). Overall, these results are consistent with the linear effect that has been previously observed in English. However, they suggest that the impact of the surprisal of the previous word varies between languages.